# The morphology and interface structure of titanium on graphene

### v1

@@ III Effect of deposition temperature on Ti morphology
As a means of Ti morphology control, we now discuss samples where Ti has been deposited on Gr at elevated temperatures. Figure 3(a) shows a TEM image of 1L and 2L Gr after 1 nm Ti deposition while heating to 400 °$\degree$°C. Additional images of 1L, 2L, and 3L Gr with Ti deposited at room temperature, 300 °$\degree$°C, and 400 °$\degree$°C are provided in Fig. S4 and S5.
There is a marked difference in the Ti morphology between 1L and 2L or 3L Gr after a heated Ti deposition. The difference is more evident than that seen for room temperature deposition. On 2L and 3L Gr heated deposition produces faceted islands that approach the Winterbottom equilibrium shape of hexagonally close-packed Ti (inset, Fig. 3(a); see Methods for calculation details) [58]. It is typical for metallic islands to exhibit dendritic growth at lower temperatures and compact, faceted growth at higher temperatures, due to increased adatom diffusion along island edges as well as diminished kinetic limitations [1]. The diffraction pattern for 2L Gr (Fig. 3(b)) shows that the faceted Ti islands are epitaxially aligned with the Gr. Furthermore, we estimate that 15-20% of the islands are rotated 30°$\degree$ with respect to the expected epitaxial alignment. Hence, the effect of increasing the temperature is generally to form compact and faceted islands. This is similar to literature observations of Nb, Au, and Ti on few-layer Gr [40], corresponding to a reduction of the influence of kinetic factors at elevated temperatures.
@@
Figure 3: Layer and substrate dependent Ti morphology after depositing Ti at 400 °$\degree$°C. (a) TEM image of a region with 1L and 2L Gr after Ti deposition at 400 °$\degree$°C. The inset shows the calculated equilibrium Winterbottom shape of Ti which is bounded by the top (0001) surface (red) and the $\{11\bar{2}1\}$ side facets (gold). (b) Diffraction patterns from the 1L and 2L region. Gr and Ti reflections are indexed with black and green font, respectively. The black and green arrow point to a Gr $\{2\bar{1}\bar{1}0\}$ and a second order Ti $\{10\bar{1}0\}$ spot.
@@
On 1L Gr, however, we find that heated deposition produces irregularly shaped islands. This can be understood if reactions with Gr hinder the free movement of the incoming Ti atoms. We have in several instances observed visible damage to the 1L Gr after heated Ti deposition, such as holes or rolled over Gr (see top right corner of the image in Fig. 3(a)), indicating damage to 1L Gr during heated deposition. This is also consistent with Raman spectroscopy data, presented below, showing more defect formation and/or stronger Ti-Gr interaction on 1L Gr. Thus, these results indicate that the suspended 1L Gr becomes damaged when depositing Ti while heating the Gr to 300-400 °$\degree$°C.
@@
Figure 4: The effect of substrate on Ti morphology. (a, b) Depositions at room temperature. AFM scans of 1.1 nm Ti deposited onto regions with (a) suspended, (b) an area containing both suspended and SiN-supported 1L Gr as indicated with arrows. (c, d) Depositions at 300 °$\degree$°C. DF-STEM images of (c) 1L Gr and (d) 2L Gr.
@@
In summary, we find that the result of high temperature deposition depends strongly on the Gr thickness. For 2L and 3L Gr, depositing Ti at 300 or 400 °$\degree$°C results in the formation of faceted and compact islands with morphologies close to the Winterbottom shape, as expected when thermodynamic factors determine island shape. On 1L Gr, depositions at 300 or 400 °$\degree$°C lead to irregularly shaped islands which can be explained by strong interactions with the Gr and defect formation.
IV Substrate dependence of Ti morphology and nucleation characteristics
@@
V Defect formation in Gr during Ti deposition
@@
To evaluate defect formation in the Gr during Ti evaporation we obtained Raman spectroscopy data from samples with 1L, 2L, and 3L suspended Gr after Ti deposition at room temperature (Fig. 5(a)). Raman spectra of Gr show characteristic G and 2D peaks at $\sim$1580 cm-1 and 2690 cm-1, respectively. Literature indicates that the D-peak at $\sim$1350 cm-1 appears only in the presence of certain defects. The D-peak is associated with vacancy-type and sp3-type defects such as TiC bonds [63], while charged impurities, intercalants and strain do not produce a D-peak [64]. In our case, suspended 1L Gr with Ti deposited at room temperature exhibits a D-peak, while 2L and 3L Gr do not show a noticeable D-peak. This suggests that defect formation, such as vacancy-type and sp3-type defects, is strongly reduced or even eliminated for 2L and 3L Gr. Similarly, Ti deposition at 400 °$\degree$°C leads to a D-peak on 1L Gr, which is relatively larger than that from 1L Gr room temperature deposition.
@@
The defect concentration is quantified using the intensity ratio between the D- and the G-peak, $I_{\mathrm{D/G}}$ [65]. For 1L Gr at room temperature ($I_{\mathrm{D/G}}$=1.3), the defect density is 1.4$\cdot$104 µm-2. With a Ti nucleation density of 960 µm-2 on 1L suspended Gr at room temperature this suggests that each Ti island contains $\sim$15 point defects in the underlying 1L Gr support. For 1L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}$=2.9), the defect density increases to $5.4\cdot 10^{4}$ µm-2. For 2L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}\sim$0.1), we estimate an upper bound of $<1.6\cdot 10^{3}$ µm-2 defects. This upper bound is based on measurements in Ref. [65], which considers a $I_{\mathrm{D/G}}$-ratio down to 0.2, corresponding to this defect density. These results indicate a strong Gr thickness and temperature dependence on the defect formation in Gr after Ti deposition.
@@
If we compare these Raman spectroscopy data with the structural characterization from Fig. 1 and 3, we obtain a more complete picture of the defect structure of the samples. For room temperature deposition, 1L Gr exhibits many more defects in Raman spectroscopy than 2L or 3L Gr, although structural changes (nucleation density and island size) change more gradually with Gr layer number. In general, high-temperature depositions show more defects compared to room temperature depositions. For high temperature depositions, 1L Gr shows at least an order of magnitude higher defect density than 2L or 3L Gr. For the high temperature case, the structural changes are also dramatic: 1L Gr displays irregularly shaped Ti islands and visible holes, whereas 2L and 3L Gr show compact and faceted islands. Our previous analysis shows that increasing surface roughness results in a stronger Ti-Gr interaction. Since surface roughness of free-standing Gr decreases with increasing thickness [54], our Raman and structural TEM characterization align well with previous studies linking the reactivity of Gr to its surface roughness [66, 67, 68].
VI Conclusions
@@
To conclude, we report a complicated dependence of the Ti-Gr interface morphology and defect formation on layer thickness, substrate, and temperature. We summarize our findings in Table 2. In general, Ti nucleation density depends on the Gr thickness. The island shape and degree of faceting depend on the temperature, in a way expected from kinetic considerations. We found that the Gr thickness dependence on Ti diffusivity and defect formation as measured by Raman spectroscopy depends on surface roughness, where increased surface roughness leads to a stronger Ti-Gr interaction. This was supported by DFT simulations which do not predict a strong Gr thickness dependence on the interaction with Ti for pristine Gr, suggesting that extrinsic factors such as roughness are important.
@@
Results for 1L Gr stand out from the results of 2L or 3L Gr. 1L Gr displays at least an order magnitude larger defect density and completely different Ti island morphologies as compared to thicker Gr flakes. We believe these differences arise due to structural and mechanical differences between 1L and thicker Gr flakes, where 1L Gr is more flexible, exhibits larger surface roughness and, therefore, displays lower Ti diffusivity, and is more easily damaged.
@@
These results highlight the extreme sensitivity of processing conditions on Ti deposition on Gr and emphasize the importance of considering factors such as the Gr thickness, the substrate on which Gr is placed, surface roughness, and strain for applications where the microscopic nature of the metal-Gr interface is critical to performance. This is in addition to the usual factors considered in metal nucleation and growth, such as deposition rate and substrate temperature.
@@
The summary in Table 2 offers guidelines for creating Ti-Gr interfaces with ideal properties. For large single crystals with minimized defect formation, it would be interesting to investigate Ti growth on Gr supported by hexagonal boron nitride or another thicker 2D material, as this should suppress damage formation, allowing the growth of facetted islands, while also providing a flat substrate minimizing nucleation density and creating large islands. It may be further possible to grow large, faceted islands by increasing the temperature. Also, interesting would be extending this work to other 2D materials like transition metal dichalcogenides (TMDs), where the Ti-TMD interface is heavily utilized in nanoelectronics research and control of the reactions at the interface are desired [12, 13]. Future studies should also examine the properties of Ti as an adhesion layer to other metals, and how the properties of other metals are affected by the Ti layer. It may be possible to grow a metallic overlayer which is epitaxial with respect to the epitaxial Ti film, and therefore create epitaxial multilayers on Gr. Furthermore, the faceted Ti islands obtained by evaporating Ti at elevated temperatures may be useful for optoelectronic applications and may serve as seeds for other metals on Gr to create complicated multilayer island structures.
VII Methods
VII.1 Sample Fabrication
@@
Gr was obtained by exfoliating natural graphite (NaturGrafit GmbH, Germany) onto oxygen plasma treated substrates of 90 nm thick SiO2 on Si using 3M Magic Scotch tape. We identified suitable flakes by their contrast using optical microscopy as shown in Fig. S1. We then transferred crystals to either homemade TEM compatible sample carriers or ”location-tagged TEM grids” purchased from Norcada, Inc., Canada. The homemade grids consist of 300 nm thick SiN membrane windows with multiple $\sim$3 µm diameter holes etched in the window. The Norcada TEM grids have a 200 nm thick SiN membrane with several 2 µm holes. The Gr is transferred using wedging transfer, based on cellulose acetate butyrate as polymer handle for the transfer, which we have previously shown produces ultra-clean suspended Gr after thermal annealing in vacuum [10, 51]. After the transfer we bake the TEM sample carriers to 140 °$\degree$°C to improve adhesion. Then we dissolve the polymer in acetone and rinse the TEM sample carriers in isopropanol. Finally, we dry the sample with critical point drying to avoid damage to the suspended Gr sample.
VII.2 Ti deposition
@@
For the suspended Gr, to remove polymeric residue, we anneal the sample overnight at 500 °$\degree$°C prior to evaporation. The annealing takes place in an ultrahigh vacuum chamber that also contains the Ti evaporator. We deposit Ti using an electron-beam evaporator at a typical rate of around 0.3 Å/min and a thickness of approximately 1 nm.
VII.3 Electron microscopy
@@
VII.4 Data analysis
Ti island nucleation density was estimated by delineating Ti islands by hand and measuring their areas using imageJ [10]. Fig. S2 and S3 shows examples of these measurements.
VII.5 Raman spectroscopy
Raman spectroscopy was performed using a Renishaw inVia confocal Raman microscope with a 532 nm excitation laser. We acquired maps with 0.5 µm spacing between spectra. We then sum 10 spectra to obtain the data shown in Fig. 5.
VII.6 Atomic force microscopy

### v2

@@ I.2 Effect of deposition temperature on Ti morphology
As a means of Ti morphology control, we now discuss samples where Ti has been deposited on Gr at elevated temperatures. Figure 3(a) shows a TEM image of 1L and 2L Gr after 1 nm Ti deposition while heating to 400 °$\degree$°C. Additional images of 1L, 2L, and 3L Gr with Ti deposited at room temperature, 300 °$\degree$°C, and 400 °$\degree$°C, corresponding to 0.30-0.35 $T_{\mathrm{M}}$, are provided in Fig. S4-S8.
@@
Figure 3: Layer dependent Ti morphology after depositing Ti at 400 °$\degree$°C. (a) TEM image of a region with 1L and 2L Gr after Ti deposition at 400 °$\degree$°C. The inset shows the calculated equilibrium Winterbottom shape of Ti which is bounded by the top (0001) surface (red) and the $\{11\bar{2}1\}$ side facets (gold). (b, c) Diffraction patterns from the (b) 1L and (c) 2L region. Gr and Ti reflections are indexed with black and green font, respectively. The green arrow points to a second order Ti $\{10\bar{1}0\}$ spot. (d) TEM image of a different region on the same sample with 2L Gr. The linear feature is likely a fold in the Gr (see also Fig. S6, caption). (e, f) Diffraction patterns from the regions indicated with the blue and red circle in (d).
There is a marked difference in the Ti morphology between 1L and 2L or 3L Gr after a heated Ti deposition. The difference is more evident than that seen for room temperature depositions. On 2L and 3L Gr, heated deposition produces faceted islands that approach the Winterbottom equilibrium shape of hexagonally close-packed Ti (inset, Fig. 3(a); see Methods for calculation details) [60]. It is typical for metallic islands to exhibit dendritic growth at lower temperatures and compact, faceted growth at higher temperatures, due to increased adatom diffusion along island edges as well as diminished kinetic limitations [1]. Our islands in fact exhibit a variety of morphologies, similar to the coexistence between different morphologies that is known for Pt and Al islands deposited at intermediate temperatures between 0.2-0.3 $T_{\mathrm{M}}$ (see Fig. 3.15 in [2]). We expect the morphology to converge towards the Winterbottom shape at temperatures above 0.3 $T_{\mathrm{M}}$: literature observations of Nb, Au, and Ti on few-layer Gr [42] also approached the Winterbottom morphology as the temperature was increased.
[…]
The diffraction pattern for 2L Gr (Fig. 3(b)) shows that the faceted Ti islands are epitaxially aligned with the Gr, although we find that for Ti deposited at 400 °$\degree$°C, around 20% of the islands are rotated 30°$\degree$ with respect to the expected epitaxial alignment. Figure S7 shows several images from the sample shown in Fig. 3 used to obtain statistics of the island rotations. Figure 3(d) highlights one overview TEM image from a 2L Gr region of this sample. The blue and red dotted circles indicate the location of the selected area aperture used to acquire the diffraction patterns shown in Fig. 3(e) and (f), respectively. The red dotted circle encloses a Ti island with a 30°$\degree$ rotation. We believe that thermodynamic barriers to forming these less favorable configurations may be overcome by thermal energy at higher temperature. Since large scale defects are not observed in the rotated islands, we expect the rotated interface structure may be present at nucleation, and be preserved as the islands grow larger.
@@
On 1L Gr we observe a different behavior compared to 2L and 3L Gr. In this case, heated deposition produces irregularly shaped islands. This can be understood if reactions with Gr hinder the free movement of the incoming Ti atoms. We have in several instances observed visible damage to the 1L Gr after heated Ti deposition, such as holes or rolled over Gr (see top right corner of the image in Fig. 3(a) and Fig. S8 for higher-magnification images), indicating damage to 1L Gr during heated deposition. This is also consistent with Raman spectroscopy data, presented below, showing more defect formation and/or stronger Ti-Gr interaction on 1L Gr. Thus, these results indicate that the suspended 1L Gr becomes damaged when depositing Ti on Gr while heated to 300-400 °$\degree$°C.
@@
Figure 4: The effect of substrate on Ti morphology. (a, b) Ti depositions at room temperature. AFM scans of 1.1 nm Ti deposited onto regions with (a) suspended, (b) an area containing both suspended and SiN-supported 1L Gr as indicated with a white arrow. (c, d) Ti depositions at 300 °$\degree$°C. DF-STEM images of (c) 1L Gr and (d) 2L Gr.
@@
In summary, we find that the result of high temperature deposition depends strongly on the Gr thickness. For 2L and 3L Gr, depositing Ti at 300-400 °$\degree$°C results in the formation of faceted and compact islands with morphologies close to the Winterbottom shape, as expected when thermodynamic factors determine island shape. Furthermore, $N$ decreases with temperature as expected for a thermally driven diffusion process, indicating that it is not governed by a few low-energy defect sites on the Gr. In contrast, on 1L Gr, depositions at 300 °$\degree$°C or 400 °$\degree$°C lead to irregularly shaped islands which can be explained by strong interactions with the Gr and defect formation.
I.3 Substrate dependence of Ti morphology and nucleation characteristics
@@
I.4 Defect formation in Gr during Ti deposition
@@
To evaluate defect formation in the Gr during Ti evaporation we obtained Raman spectroscopy data from samples with 1L, 2L, and 3L suspended Gr after Ti deposition at room temperature and 400 °$\degree$°C (Fig. 5). Raman spectra of Gr show characteristic G and 2D peaks at $\sim$1580 cm-1 and 2690 cm-1, respectively. Literature indicates that the D-peak at $\sim$1350 cm-1 appears only in the presence of certain defects. The D-peak is associated with vacancy-type and sp3-type defects, which occur when Gr forms bonds to other materials or compounds, such as TiC bonds [65]. Charged impurities, intercalants and strain do not produce a D-peak [66]. In Gr devices, the intensity of the D-peak, typically measured as the intensity ratio between the D- and the G-peak, $I_{\mathrm{D/G}}$, has been found to scale inversely with the charge carrier mobility [67, 68]. This underscores the importance of strategies for avoiding defect formation during processing when Gr is present.
In our case, suspended 1L Gr with Ti deposited at room temperature exhibits a D-peak, while 2L and 3L Gr do not show a noticeable D-peak. This suggests that defect formation, such as vacancy-type and sp3-type defects, is strongly reduced or even eliminated for 2L and 3L Gr. Similarly, Ti deposition at 400 °$\degree$°C leads to a D-peak on 1L Gr, which is relatively larger than that from the room temperature deposition.
@@
The defect concentration is quantified using the intensity ratio between the D- and the G-peak, $I_{\mathrm{D/G}}$ [69]. For 1L Gr at room temperature ($I_{\mathrm{D/G}}$=1.3), the defect density is 1.4$\cdot$104 µm-2. With a Ti nucleation density of 1280 µm-2 on 1L suspended Gr at room temperature this suggests that each Ti island contains $\sim$11 point defects in the underlying 1L Gr support. For 1L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}$=2.9), the defect density increases to $5.4\cdot 10^{4}$ µm-2. For 2L Gr at 400 °$\degree$°C ($I_{\mathrm{D/G}}\sim$0.1), we estimate an upper bound of $<1.6\cdot 10^{3}$ µm-2 defects. This upper bound is based on measurements in Ref. [69], which considers a $I_{\mathrm{D/G}}$-ratio down to 0.2, corresponding to this defect density. These results indicate a strong Gr thickness and temperature dependence on the defect formation in Gr after Ti deposition.
@@
If we compare these Raman spectroscopy data with the structural characterization from Fig. 1 and 3, we obtain a more complete picture of the defect structure of the samples. For room temperature depositions, 1L Gr exhibits many more defects in Raman spectroscopy than 2L or 3L Gr, although structural changes (nucleation density and island size) change more gradually with Gr layer number. In general, high-temperature depositions show more defects compared to room temperature depositions. For high temperature depositions, 1L Gr shows at least an order of magnitude higher defect density than 2L or 3L Gr. For the high temperature case, the structural changes are also dramatic: 1L Gr displays irregularly shaped Ti islands and visible holes, whereas 2L and 3L Gr show compact and faceted islands.
Our previous analysis showed that increasing surface roughness results in a stronger Ti-Gr interaction, and we believe the differences in defect formation for 1L-3L Gr are also related to surface roughness. The reactivity of Gr has been shown to increase with increasing surface roughness [70, 71, 72]. Since surface roughness of free-standing Gr decreases with increasing thickness [56], our Raman and structural TEM characterization align well with these previous studies. These results also suggest that placing 1L Gr on a flat substrate like hexagonal boron nitride or another relevant 2D material could be an avenue towards reducing defect formation during Ti layer growth.
II Conclusions
@@
To conclude, we report a complicated dependence of Ti-Gr interface morphology and defect formation on Gr layer thickness, substrate, and temperature, and we discuss how these can be understood through the central role of Gr roughness. Our findings are summarized in Table 2. Ti island nulceation density, shape and degree of faceting depend on the temperature in a way expected from kinetic considerations, but these parameters also depend on Gr thickness. Since DFT simulations do not predict a strong dependence of the Gr-Ti interaction on Gr thickness, we conclude that an extrinsic factor, roughness, dominates the island growth process.
@@
Results for 1L Gr stand out from the results of 2L or 3L Gr. 1L Gr displays at least an order magnitude larger defect density and completely different Ti island morphologies as compared to thicker Gr flakes. This supports the idea of these differences arising due to structural and mechanical differences between 1L and thicker Gr flakes, where 1L Gr is more flexible, exhibits larger surface roughness and, therefore, displays lower Ti diffusivity and is more easily damaged.
@@
These results highlight the extreme sensitivity of processing conditions on Ti deposition on Gr and emphasize the importance of considering factors such as the Gr thickness, the substrate on which Gr is placed, particularly its roughness, for applications where the microscopic nature of the metal-Gr interface is critical to performance. This is in addition to the usual factors considered in metal nucleation and growth, such as deposition rate and substrate temperature.
@@
The summary in Table 2 offers guidelines for creating Ti-Gr interfaces with ideal properties. For large single crystals with minimized defect formation, it would be interesting to investigate Ti growth on Gr supported by hexagonal boron nitride or another thicker 2D material, as this should suppress damage formation, allowing the growth of faceted islands, while also providing a flat substrate minimizing nucleation density and creating large islands. It may be further possible to grow large, faceted islands by increasing the temperature. Another option is a two-step process with a low-temperature initial deposition to limit interfacial defect formation, followed by a higher temperature deposition to increase Ti grain size. Additionally, it would be interesting to extend this work to other 2D materials like transition metal dichalcogenides (TMDs), where the Ti-TMD interface is heavily utilized in nanoelectronics research and control of the reactions at the interface are desired [12, 13]. Future studies should also examine the properties of Ti as an adhesion layer to other metals, and how the properties of other metals are affected by the Ti layer. It may be possible to grow a metallic overlayer which is epitaxial with respect to the epitaxial Ti film, and therefore create epitaxial multilayers on Gr. Furthermore, the faceted Ti islands obtained by evaporating Ti at elevated temperatures may be useful for optoelectronic applications and may serve as seeds for other metals on Gr to create complicated multilayer island structures.
III Methods/Experimental
III.1 Sample Fabrication
@@
Gr was obtained by exfoliating natural graphite (NaturGrafit GmbH, Germany) onto oxygen plasma treated substrates of 90 nm thick SiO2 on Si using 3M Magic Scotch tape. We identified suitable flakes by their contrast using optical microscopy as shown in Fig. S1. We then transferred crystals to either homemade TEM compatible sample carriers or ”location-tagged TEM grids” purchased from Norcada, Inc., Canada. The homemade grids consist of 300 nm thick SiN membrane windows with multiple $\sim$3 µm diameter holes etched in the window. The Norcada TEM grids have a 200 nm thick SiN membrane with several 2 µm holes. The Gr is transferred using wedging transfer, based on cellulose acetate butyrate as polymer handle for the transfer, which we have previously shown produces ultra-clean suspended Gr after thermal annealing in vacuum [10, 53]. After the transfer we bake the TEM sample carriers to 140 °$\degree$°C to improve adhesion. Then we dissolve the polymer in acetone and rinse the TEM sample carriers in isopropanol. When drying the samples we use critical point drying. This is typically used for fragile samples where surface tension could otherwise cause damage, including free-standing Gr samples [10, 42, 53, 73, 74]. Damage to Gr during a regular drying process, where a solvent like isopropanol evaporates from the surface of the TEM grid, typically results in easily observed (through optical microscopy) macroscopic effects such as torn or rolled Gr, especially in 1L Gr.
III.2 Ti deposition
@@
For the suspended Gr, to remove polymeric residue, we anneal the sample overnight at 500 °$\degree$°C prior to evaporation. The annealing takes place in an ultrahigh vacuum chamber that also contains the Ti evaporator. The pressure during the evaporation was typically between $1\times 10^{-8}$$2\times 10^{-8}$ Torr. We deposit Ti using an electron-beam evaporator at a typical rate of around 0.3 Å/min and a thickness of approximately 1 nm. Fig. S10 shows Raman spectroscopy data obtained after Gr transfer to the TEM grids (”as prepared”), after the annealing step in our UHV system (”after UHV anneal”), and after Ti deposition at 400 °$\degree$°C. This data shows that Gr transfer and UHV annealing does not lead to any detectable defect formation in the Gr, and that a D-peak in 1L Gr only appears after Ti deposition.
III.3 Electron microscopy
@@
III.4 Data analysis
Ti island nucleation density was estimated by delineating Ti islands by hand and measuring their areas using imageJ [10]. Fig. S3 and S6 show these measurements. Extracting the nucleation density for our depositions at elevated temperatures is challenging since at least 50% of islands have coalesced.
III.5 Raman spectroscopy
Raman spectroscopy was performed using a Renishaw inVia confocal Raman microscope with a 532 nm excitation laser. We acquired maps with 0.5 µm spacing between spectra. We then sum 10 spectra to obtain the data shown in Fig. 5. The laser power was set to 2.5 mW. This is well below the power required to visibly damage the Gr.
III.6 Atomic force microscopy